OpenStudy (anonymous):

2r(2r-1)=3r^2-3r+20 factoring

OpenStudy (campbell_st):

well if you distribute the left hand side, what do you get...?

OpenStudy (anonymous):

Um 3r(r-1)+20

OpenStudy (campbell_st):

so you have \[2r(2r -1) = 2r \times 2r - 2r \times 1\] can you simplify this